OpenStudy (anonymous):

How would you simplify this complex fraction: (b^(-2))-(x^(/2) / (b^(-1))+(x^(-1))

OpenStudy (anonymous):

i think the power of x is x^-2 not x^/2

OpenStudy (anonymous):

am i right

OpenStudy (anonymous):

Yeah;; That's what I meant.

myininaya (myininaya):

\[\frac{\frac{1}{b^2}-\frac{1}{x^2}}{\frac{1}{b}+\frac{1}{x}}\]

myininaya (myininaya):

\[\frac{(\frac{1}{b}-\frac{1}{x})(\frac{1}{b}+\frac{1}{x})}{\frac{1}{b}+\frac{1}{x}}\]

myininaya (myininaya):

\[=\frac{1}{b}-\frac{1}{x}\]

OpenStudy (anonymous):

Thank you very much <3

myininaya (myininaya):

:)